# FLM-Audio: Natural Monologues Improves Native Full-Duplex Chatbots via Dual Training

### v2

@@ Abstract
Full-duplex dialog models aim to listen and speak simultaneously, delivering rapid responses to dynamic user input. Among different solutions to full-duplexity, a native solution merges multiple channels in each time step, achieving the lowest latency. However, prevailing designs break down the textual monologue sentences for word-level alignment with audio streams, which degrades language modeling abilities. To help address this issue, we introduce “natural monologues”, which are composed by continuous sentences and “waiting” intervals, mimicking humanoid cognitive behavior in dialogs. We find a proper training paradigm to be critical for semantically aligning natural monologues with audio. To this end, we develop a “dual” training paradigm that alternates the position of the monologues, either leading or trailing the audio, across different training stages. A combination of our natural monologue and dual training strategy is applied in developing FLM-Audio, our 7B spoken dialog chatbot with native full-duplexity. As confirmed by experimental results, FLM-Audio achieves superior response qualities and chatting experiences while requiring significantly less training data.
22footnotetext: Indicates equal contribution.11footnotetext: Corresponding author.
@@ 1 Introduction
Humanoid responsiveness is increasingly regarded as a key capability for applied AI systems. Humans naturally respond to rapidly changing visual, auditory, and other sensory inputs with real-time speech, monologues, gestures, and actions. Achieving comparable responsiveness is therefore recognized as a critical requirement for advanced AI, particularly for higher levels of embodied intelligence such as L3+ embodied AGI towardembodied .In this paper, we focus on the audio and textual modalities, investigating humanoid responsiveness within Spoken Dialog Models (SDMs). Such responsiveness involves both humanoid dialog behaviors (e.g., natural speech style, turn-taking, and graceful handling of interruptions) and humanoid response latency (e.g., reacting promptly to dynamic environmental inputs). A common architectural principle underlying these behaviors in SDMs is the implementation of full-duplex mechanisms sdm-bench ; beyond ; full-dup .
Two major strategies have emerged for enabling full duplexity: Time-Division Multiplexing (TDM) and Native Full-duplexity (Figure 1). TDM, widely adopted in state-of-the-art audio-language models audiolm ; glm-voice ; mini-omni ; qwen-audio ; freezeomni , interleaves listening and speaking tokens within a single sequence. In each forward pass, a TDM model’s context is a concatenated stream from all input and output channels (e.g., listening, monologue text, and speaking). As the Transformer attention mechanism vaswani2017attention has a computational complexity of $O(n^{2})$, TDM significantly hampers responsiveness, resulting in full-duplex delays of up to 2 seconds beyond , and limits audio generation length to roughly 45 seconds minicpm-o . These bottlenecks become increasingly restrictive as the foundation models continue to scale up chinchilla ; mu-scaling ; o1 .
Figure 1: TDM vs. Native Full Duplexity for humanoid responsiveness.
@@
In this paper, we follow the Native Full-duplexity paradigm for its superior scaling potential, but instead introduce continuous monologue tokens, which we term “natural” monologues. Instead of aligning every token to its precise pronunciation, we generate uninterrupted token sequences in the text channel (e.g., a full sentence or paragraph) while the speaking channel concurrently produces audio. Typically, regarding aligned time steps, the textual sentence finishes much earlier than the speaking channel. During this gap, the model emits continuous <wait> tokens until the next monologue sentence is triggered. This approach preserves the language modeling strength of foundation models. Furthermore, during pre-training, transcripts and audio only need to be aligned at the sentence level, which both lowers pre-processing cost and mitigates error propagation from misaligned word timestamps. Figure 2 illustrates the contrast between alignment strategies.
@@
Incorporating natural monologues in native full-duplex paradigm is a non-trivial problem: compared to word-level alignment, a model with natural monologues needs to learn to generate text and audio simultaneously, even when their semantic contents are asynchronous (e.g., the speech channel may still be pronouncing word A while the text channel has already advanced to words B or C). Our experiments show that the optimal stream arrangement, training objective, and configuration details differ substantially from those in related work moshi . To this end, we design a “dual-phase” training scheme, where the natural monologue alternately leads or lags behind the audio channel across training stages, effectively covering both ASR- and TTS-like modes. We observe that such training strategy enables the model to handle the asynchronous semantics across long paragraphs, yielding coherent natural monologues together with human-like speech.
The contribution of this paper includes:
  • We propose a novel framework for native full-duplex audio chatbots, featuring a stream organization method based on natural monologues, as well as the corresponding complete training pipeline.
  • We release FLM-Audio, an open-source full-duplex audio-language model  [^1] https://huggingface.co/CofeAI/FLM-Audio, along with the codes for the inference and interaction pipeline  [^2] https://github.com/cofe-ai/flm-audio.
@@
  • Experimental results show that FLM-Audio outperforms native full-duplex baselines with much less post-training data, and surpasses state-of-the-art models in humanoid responsiveness tests including automatic and human evaluation.
@@
Table 1: Summary of related work. Full-Duplex stands for whether the model demonstrates capabilities to process input and output simultaneously, with the minimal requirement of reacting promptly to interruptions in the listening channel. E2E denotes whether the model is end-to-end, especially, an E2E model learns to directly generate audio tokens instead of relying on external ASR/TTS modules (though external token-to-wave audio decoders may still be used). Following sdm-bench , we also summarize whether the full-duplex speech-to-speech pipeline is open-sourced (S2S Release).
@@
MiniCPM-Duplex (beyond, ) TDM en
MiniCPM-Duo (cpm-duo, ) CDM en
MinMo (minmo, ) TDM multi
GLM-4-voice (glm-voice, ) - - en,zh
Kimi-Audio (kimi-audio, ) - - en,zh
Freeze-Omni (freezeomni, ) TDM en,zh
OmniFlatten (omniflatten, ) TDM en,zh
Moshi (moshi, ) Native en
@@ 2 FLM-Audio: Model Design
In this section, we introduce FLM-Audio, a native full-duplex model utilizing natural monologues through multi-stage training. FLM-Audio follows the Native approach (Figure 1, right), merging listening, speaking, and monologue channels at each autoregressive (AR) step of the backbone model. As discussed above, this approach avoids time-slice sharing by time-division multiplexing (TDM). We summarize previous work in Table 1, observing that most existing audio-language models (as well as other omnimodal visual-language models such as MiniCPM-o minicpm-o and Qwen2.5-Omni qwen-omni ) use TDM as a solution for full duplexity, with Moshi moshi being a notable exception. While FLM-Audio adopts a similar backbone design to Moshi, we introduce key differences and improvements in stream organization, text–audio alignment, and the training pipeline.
@@ 2.1 Backbone Structure
Due to limitations in computational resources, we restrict the scale of our foundation model to $\sim$7B rather than using larger models such as Tele-FLM-52B teleflm . Since our goal is to support both English and Chinese, we also exclude English-only model families such as Llama llama3 and Mistral mistral . Accordingly, we adopt a 7B-parameter autoregressive LLM as the backbone, initialized from the language model component of Qwen-2.5-VL qwen2.5vl  [^3] We select Qwen-2.5-VL instead of a unimodal LLM to retain visual understanding capability for program management purposes..
We follow the RQ-Transformer architecture uniaudio ; hierarchical employed by Moshi moshi for streaming audio processing. This choice ensures better comparability, and we believe that in LLM-driven research, meaningful gains can stem directly from innovations in data organization, alignment strategies, and training paradigms, even when the core architecture is kept intact. Audio waveforms are discretized at 12.5 frames per second, with 8 audio tokens per frame. In each time step of the backbone model, a depth transformer moshi ; uniaudio ; hierarchical takes the last-layer hidden states from the backbone model as input, and generates 8 audio tokens (1 semantic tokens followed by 7 acoustic tokens) in a locally autoregressive manner. Streaming Mimi encoders and decoders  [^4] https://huggingface.co/kyutai/mimi serve as bridges between tokens and waveforms.
Formally, with $e$ denoting embeddings of textual or audio tokens, the backbone model $F$ is defined as:
@@
We observe that the hidden state $h_{t}$ is sufficiently informative for textual, semantic, and acoustic generation. As a result, the depth Transformer can depend solely on the local $h_{t}$, without the need to re-aggregate $O(N^{2})$ contextual information as required in the “talker”-like architectures employed in other related work qwen-omni .
@@
2.2 Natural Monologues
Even within a single utterance, textual and audio tokens are inherently asynchronous: one second of speech— represented by 12.5 frames of audio features—typically corresponds to only 3–4 monologue tokens. To address this mismatch, Moshi moshi adopts a token-level alignment strategy, where textual tokens are split with special <pad> and <end-of-pad> tokens, ensuring each token to appear precisely at the time it is spoken (Figure 2, left). While effective, this approach has two major drawbacks: (1) it requires fine-grained, word-level timestamps for training annotations, which significantly increases data processing cost and introduces vulnerability to cascading alignment errors; and (2) it diverges from humanoid dialog patterns. In natural conversations, humans think, listen, and speak concurrently, with internal monologues forming a coherent, forward-moving stream that generally precedes speech. From an empirical perspective, fragmenting sentences into isolated word-level tokens undermines the language modeling capacity of the backbone, as noticed in the original work moshi . Consistently, related work has also reported limited instruct-following performance for Moshi minmo ; omniflatten ; sdm-bench .
To overcome these limitations, FLM-Audio adopts a “Natural Monologues” strategy, Instead of aligning text and audio at the word level, monologues are aligned at the sentence level, represented as continuous token sequences interleaved with continuous <wait> token sequences between sentences or paragraphs. The natural monologues can either lead or follow the spoken audio.
@@ 2.2 Natural Monologues
Natural monologues requires only sentence-level transcripts for training, which drastically reduces annotation cost and scales more easily to large datasets. Furthermore, it preserves the autoregressive language modeling capabilities of the pretrained backbone, supporting both natural dialog generation and responsive full-duplex speech.
@@ 3 FLM-Audio: Dual Training Paradigm
Although both FLM-Audio and Moshi moshi adopt a RQ-Transformer uniaudio ; hierarchical model architecture, Moshi’s training pipeline can not be trivially transferred to FLM-Audio. This is due to fundamental differences in monologue alignment strategies, as discussed in Section 2.2. We summarize the distinctions across post-training and fine-tuning stages in Table 2.
Both models undergo four stages in total, including two post-training and two fine-tuning stages. However, to better exploit the language modeling benefits of natural monologues, FLM-Audio features special designs to enhance sentence-level alignment with both listen and speak channels during the early stages. Because our framework incorporates both TTS-style and ASR-style data formats throughout post-training and fine-tuning, we term this approach the “Dual Training Paradigm”.
@@ 3.1 Stage-1: Post-training
The objective of post-training is to equip the pretrained language model with both listening and speaking abilities. At this stage, large-scale audio-text data is used to jointly train autoregressive modeling of acoustic codes and semantic alignment between textual and audio modalities. For data processing, we compile a corpus of roughly 1 million hours of speech audio covering multiple Chinese and English sources including audio books, podcasts, TV shows, vlogs, etc. The audios are transcribed by FunASR funasr for Chinese and Whisper whisper for English, followed by text filtering to remove erroneous or noisy samples.
The post-training has two sub-stages: in the first sub-stage (Post-training-1), the entire transcribed corpus is used as training data; in the second sub-stage (Post-training-2), we further incorporate a suite of open-sourced, human-annotated ASR datasets including ST_CMDS [^5] https://openslr.org/38/, Aishell3 aishell3 , Magicdata [^6] https://www.openslr.org/68/, primewords_md_2018_set1  [^7] https://www.openslr.org/47/, and Thchs30 thchs . In the second stage, to balance annotation quality, we down-sample the transcribed corpus from Post-training-1 by half (as it relies on automatic transcripts) and up-sample the human-annotated datasets by a factor of 5 to emphasize their finer-grained accuracy.
In both sub-stages, only sentence-level timestamps are extracted. Each aligned (audio clip, textual sentence) pair is tokenized and organized into two dual formats (Figure 3, left):
TTS Style. The listening channel is filled with an empty token (ID 2048 in the Mimi tokenizer). The monologue text is placed continuously on the text channel, while speech codes are aligned on the speaking channel, beginning two tokens after the start of the text. The text channel is filled with <wait> tokens until the speech output finishes. Different aligned pairs are concatenated, separated by random silence, and padded to a uniform length of 8192.
ASR Style. The speaking channel is filled with empty tokens. Speech codes are placed on the listening channel, followed by the monologue text placed continuously on the text channel starting immediately after the listening codes, effectively forming an ASR-style task.
@@
$L=\alpha_{1}*\text{CE}_{\text{speak, semantic}}+\alpha_{2}*\text{CE}_{\text{speak, acoustic}}+\beta*\text{CE}_{\text{mono}}+\gamma*\text{CE}_{\text{wait}},$ (3)
@@
This setup differs substantially from Moshi moshi : which reported optimal values of $\gamma=0.5$ for their word-level <pad> tokens, and $\alpha_{1}=100,\alpha_{2}=1$. They also leveraged text-token masking and separate optimizers for the backbone and the depth model, whereas we observed such techniques to be unnecessary for training FLM-Audio.
Our total training corpus is approximately 1 million hours—considerably smaller than Moshi’s (8+ million hours) as well as other related work minmo ; kimi-audio ; glm-voice . Nevertheless, results show that our dual training paradigm achieves comparable or superior performance (Section 4).
Table 2: Training Paradigm: FLM-Audio and Moshi.
Moshi moshi Helium 1-channel
2-channel
semi-duplex
full-duplex
dialog
full-duplex
instruct
FLM-Audio Qwen-2.5-vl
2-channel
coarse
2-channel
fine
semi-duplex
w/ ASR
full-duplex
w/o ASR

### v3

# FLM-Audio: Natural Monologues Improve Native Full-Duplex Chatbots via Dual Training

@@ Abstract
Full-duplex dialog models aim to listen and speak simultaneously, delivering rapid responses to dynamic user input. Among different solutions to full-duplexity, a native solution merges multiple channels in each time step, achieving the lowest latency. However, prevailing designs break down the textual monologue sentences for word-level alignment with audio streams, which degrades language modeling abilities. To help address this issue, we introduce “contiguous monologues”, which are composed by continuous sentences and “waiting” intervals, mimicking human-like cognitive behavior in dialogs. We find a proper training paradigm to be critical for semantically aligning contiguous monologues with audio. To this end, we develop a “dual” training paradigm that alternates the position of the monologues, either leading or trailing the audio, across different training stages. A combination of our contiguous monologue and dual training strategy is applied in developing FLM-Audio, our 7B spoken dialog chatbot with native full-duplexity. As confirmed by experimental results, FLM-Audio achieves superior response qualities and chatting experiences while requiring significantly less training data.
Machine Learning, ICML
@@ 1 Introduction
Human-like responsiveness is increasingly regarded as a key capability for applied AI systems. Human respond to rapidly-changing multimodal inputs with real-time speech, monologues, gestures, and actions. Therefore, achieving comparable responsiveness is recognized as a critical requirement for advanced AI, particularly for higher levels of embodied intelligence such as L3+ embodied AGI (Wang and Sun, 2025). In this paper, we focus on the audio and textual modalities, investigating human-like responsiveness with Spoken Dialog Models (SDMs). Such responsiveness is two-folds: it involves both human-like dialog behaviors (e.g., natural speech style, turn-taking, and graceful handling of interruptions) and human-like response latency (e.g., reacting promptly to dynamic environmental inputs). A common architectural principle underlying these behaviors is the implementation of full-duplex mechanisms (Lin et al., 2025; Zhang et al., 2024b; Wang et al., 2024a).
Two major strategies have emerged for full duplexity: Time-Division Multiplexing (TDM) and Native Full-duplexity (Figure 1). TDM, widely adopted in state-of-the-art audio-language models (Borsos et al., 2023; Zeng et al., 2024; Xie and Wu, 2024; Chu et al., 2023; Wang et al., 2024b), interleaves listening and speaking tokens within a single sequence. In each forward pass, a TDM model’s context is a concatenated stream from all input and output channels (e.g., listening, monologue text, and speaking). As the Transformer attention mechanism (Vaswani et al., 2017) has a computational complexity of $O(n^{2})$, TDM significantly hampers responsiveness, resulting in full-duplex delays of up to 2 seconds  [^1] This typically depends on the TDM chunk size, which can not be very small for semantic continuousness. (Zhang et al., 2024b), and limiting maximum generation length to roughly 45 seconds (Yuan Yao and others, 2025). These bottlenecks become increasingly restrictive as the foundation models continue to scale up (Hoffmann et al., 2022; Yao and Wang, 2023; Jaech et al., 2024).
Figure 1: TDM vs. Native Full Duplexity for human-like responsiveness.
@@
In this paper, we follow the Native Full-duplexity paradigm for its superior scaling potential, but instead introduce continuous monologue tokens, which we term “contiguous” monologues. Instead of temporally aligning every token to its pronunciation, we generate uninterrupted token sequences in the text channel (e.g., a full sentence or paragraph) while the speaking channel concurrently produces audio. Typically, the textual sentence finishes much earlier than the speaking channel due to different bit rates. During this gap, the model emits continuous ¡wait¿ tokens until the next monologue sentence is triggered. This approach preserves the language modeling strength of foundation models. Furthermore, during pre-training, transcripts and audio only need to be aligned at the sentence level, which both lowers pre-processing cost and mitigates error propagation from misaligned word timestamps. Figure 2 illustrates the contrast between alignment strategies.
@@
Incorporating contiguous monologues in native full-duplex paradigm is a non-trivial problem: compared to word-level alignment, a model with contiguous monologues needs to learn to generate text and audio simultaneously, even when their semantic contents are asynchronous (e.g., the speech channel may still be pronouncing word A while the text channel has already advanced to words B or C). Our experiments show that the optimal stream arrangement, training objective, and configuration details differ substantially from those in related work (Défossez et al., 2024). To this end, we design a “dual” training scheme, where the contiguous monologue alternately leads or lags behind the audio channel across training stages, effectively covering both ASR- and TTS-like modes. We observe that such training strategy enables the model to handle the asynchronous semantics across long paragraphs, yielding coherent contiguous monologues and human-like natural speech at the same time.
@@
The contributions of this paper include: (1) We propose a novel framework for native full-duplex audio chatbots, featuring a stream organization method based on contiguous monologues, as well as the corresponding complete training pipeline. (2) We release FLM-Audio, an open-source full-duplex audio-language model, along with the codes for the inference and interaction pipeline. Urls will be available upon publication. (3) Experimental results show that FLM-Audio outperforms native full-duplex baselines with much less post-training data, and surpasses state-of-the-art models in human-like responsiveness tests including automatic and human evaluation.
@@
Table 1: Summary of related work. Full-Duplex stands for whether the model demonstrates capabilities to listen and speak simultaneously, with the minimal requirement of reacting promptly to interruptions in the listening channel. E2E denotes whether the model is end-to-end: an E2E model learns to directly generate audio tokens instead of relying on external ASR/TTS modules (though external token-to-wave audio decoders may still be used). Following (Lin et al., 2025), we also summarize whether the full-duplex speech-to-speech pipeline is open-sourced (S2S Release).
@@
MiniCPM-Duplex (Zhang et al., 2024b) TDM en
MiniCPM-Duo (Xu et al., 2024a) CDM en
MinMo (Chen et al., 2025) TDM multi
GLM-4-voice (Zeng et al., 2024) - - en,zh
Kimi-Audio (Ding et al., 2025) - - en,zh
Freeze-Omni (Wang et al., 2024b) TDM en,zh
OmniFlatten (Zhang et al., 2024a) TDM en,zh
Moshi (Défossez et al., 2024) Native en
@@ 2 FLM-Audio: Model Design
In this section, we introduce FLM-Audio, a native full-duplex model utilizing contiguous monologues through multi-stage dual training. FLM-Audio follows the native full-duplex approach (Figure 1, right), merging listening, speaking, and monologue channels at each autoregressive (AR) step of the backbone model. As discussed above, this approach avoids time-slice sharing by time-division multiplexing (TDM). We summarize previous work in Table 1, observing that most existing audio-language models (as well as other omnimodal visual-language models such as MiniCPM-o (Yuan Yao and others, 2025) and Qwen2.5-Omni (Xu et al., 2025)) use TDM as a solution for full duplexity, with Moshi (Défossez et al., 2024) being a notable exception. While FLM-Audio adopts a similar backbone design to Moshi, we introduce key differences and improvements in stream organization, text–audio alignment, and the training pipeline.
@@ 2.1 Backbone Structure
Due to limitations in computational resources, we restrict the scale of our foundation model to $\sim$7B rather than using larger models such as Qwen-72B (Yang et al., 2025) and Tele-FLM-52B (Li et al., 2024). Since our goal is to support both English and Chinese, we also exclude English-only model families such as Llama (Meta, 2024). We opt to adopt a 7B-parameter autoregressive LLM as the backbone, initialized from the language model component of Qwen-2.5-VL (Bai et al., 2025) [^2] We choose to use Qwen-2.5-VL to retain visual capability for program management purposes..
We follow the RQ-Transformer architecture (Yang et al., 2023; Zhu et al., 2024) employed by Moshi (Défossez et al., 2024) for streaming audio processing. This choice ensures better comparability, and we believe that in LLM-driven research, meaningful gains can stem directly from innovations in data organization, alignment strategies, and training paradigms, even when the core architecture is kept intact. Audio waveforms are discretized at 12.5 frames per second, with 8 audio tokens per frame. In each time step (1 frame), a depth transformer (Défossez et al., 2024; Yang et al., 2023; Zhu et al., 2024) takes the last-layer hidden states from the backbone model as input, and generates 8 audio tokens (1 semantic tokens followed by 7 acoustic tokens) in a locally autoregressive manner. Streaming Mimi encoder and decoder  [^3] https://huggingface.co/kyutai/mimi serve as bridges between tokens and waveforms.
With $e$ denoting the embeddings of textual or audio tokens, the backbone model $F$ is defined as:
@@
We observe in experiments that the hidden state $h_{t}$ is sufficiently informative for textual, semantic, and acoustic generation. As a result, the depth Transformer can depend solely on the local $h_{t}$, without the need to re-aggregate $O(N^{2})$ contextual information as required in the “talker-like” architectures employed in other related work like Qwen2.5-omni (Xu et al., 2025).
@@
2.2 Contiguous Monologues
Even within a single utterance, textual and audio tokens are inherently asynchronous: one second of speech— represented by 12.5 frames of audio features—typically corresponds to only 3–4 monologue tokens. To address this mismatch, Moshi (Défossez et al., 2024) adopts a token-level alignment strategy, where textual tokens are split with special ¡pad¿ and ¡end-of-pad¿ tokens, ensuring each token to appear precisely at the time it is spoken (Figure 2, left). While effective, this approach has two major drawbacks: (1) it requires fine-grained, word-level timestamps for training annotations, which significantly increases data processing cost and introduces vulnerability to cascading alignment errors; and (2) it diverges from human-like dialog patterns. In natural conversations, humans think, listen, and speak concurrently, with internal monologues forming a coherent, forward-moving stream that generally precedes speech. From an empirical perspective, fragmenting sentences into isolated word-level tokens undermines the language modeling capacity of the backbone, as noticed in the original work (Défossez et al., 2024). Consistently, related work has also reported limited instruct-following performance for Moshi (Chen et al., 2025; Zhang et al., 2024a; Lin et al., 2025).
To overcome these limitations, FLM-Audio adopts a “contiguous monologues” strategy: instead of aligning text and audio at word-level, the monologues are represented as continuous token sequences, separated into sentences. Importantly, this setting mirrors human-like cognitive behaviors. The contiguous monologues can either lead or follow the spoken audio.
@@ 2.2 Contiguous Monologues
Contiguous monologues requires only sentence-level transcripts for training, which drastically reduces annotation cost. Furthermore, it preserves the autoregressive language modeling capabilities of the pretrained backbone, supporting both natural dialog generation and responsive full-duplex speech.
@@ 3 FLM-Audio: Dual Training Paradigm
Although both FLM-Audio and Moshi adopt a RQ-Transformer (Yang et al., 2023; Zhu et al., 2024) model architecture, Moshi’s training pipeline can not be trivially transferred to FLM-Audio. This is due to fundamental differences in monologue alignment strategies, as discussed in Section 2.2. Because our framework incorporates both TTS-style and ASR-style data formats throughout post-training and fine-tuning, we term this approach the “Dual Training Paradigm”. We summarize the distinctions across post-training and fine-tuning stages compared to Moshi in Appendix A.
@@ 3.1 Stage-1: Post-training
The objective of post-training is to equip the pretrained language model with both listening and speaking abilities. At this stage, large-scale audio-text data is used to train both autoregressive modeling of acoustic codes and semantic alignment between textual and audio modalities. For data processing, we compile a corpus of roughly 1 million hours of speech audio covering multiple Chinese and English sources including audio books, podcasts, TV shows, vlogs, etc. The audios are transcribed by FunASR (Gao et al., 2023) for Chinese and Whisper (Radford et al., 2023) for English, followed by text filtering to remove erroneous, harmful, or noisy samples.
The post-training has two sub-stages: in the first sub-stage (Post-training-1), the entire transcribed corpus is used as training data. In the second sub-stage (Post-training-2), we further incorporate a suite of open-sourced, human-annotated ASR datasets, including ST_CMDS [^4] https://openslr.org/38/, Aishell3 (Shi et al., 2020), Magicdata [^5] https://www.openslr.org/68/, primewords_md_2018_set1  [^6] https://www.openslr.org/47/, and Thchs30 (Wang and Zhang, 2015).
To balance annotation quality, we down-sample the transcribed corpus from Post-training-1 by half (as it relies on automatic transcripts) and up-sample the human-annotated datasets by a factor of 5 to emphasize their finer-grained accuracy. In both sub-stages, only sentence-level timestamps are extracted. Each aligned (audio clip, textual sentence) pair is tokenized and organized into two dual formats (Figure 3, left):
4 Curated open-sourced ASR datasets
TTS Style. The listening channel is filled with empty tokens with all-0 embedding. The monologue is placed continuously on the text channel, while speech codes are filled into the speaking channel, beginning two tokens after the start of the text. Different aligned pairs are concatenated, separated by random silence, and padded to a uniform length of 8192.
ASR Style. The speaking channel is filled with empty tokens. Speech codes are placed on the listening channel, followed by the monologue text, effectively forming an ASR-style task.
@@ 4 Curated open-sourced ASR datasets
$L$ $=\alpha_{1}*\text{CE}_{\text{speak, semantic}}+\alpha_{2}*\text{CE}_{\text{speak, acoustic}}$
$+\beta*\text{CE}_{\text{mono}}+\gamma*\text{CE}_{\text{wait}},$
@@
This setup differs substantially from Moshi (Défossez et al., 2024), which reported optimal values of $\gamma=0.5$ for their word-level <pad> tokens, $\alpha_{1}=100$, and $\alpha_{2}=1$. They also leveraged text-token masking and separate optimizers for the backbone and the depth model, whereas we observed such techniques to be unnecessary for training FLM-Audio.
Our stage-1 training corpus is approximately 1 million hours, considerably smaller than Moshi (8+ million hours) and other related work (Chen et al., 2025; Ding et al., 2025; Zeng et al., 2024). Nevertheless, FLM-Audio achieves comparable or superior performance in certain tasks (Section 5).